Two triangles are similar. The sides of the first triangle are 2, 4, and 6 units. The largest side of the second triangle is 24 units. Find the perimeter of the second triangle.

Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1659)

Use proportions to find the missing sides.

6/24 = 4/x {match the corresponding sides among 1st and 2nd triangles}
6x = 96 {cross-multiplied}
x = 16 {divided each side by 6}

6/24 = 2/x {matched corresponding sides}
6x = 48 {cross-multiplied}
x = 8 {divided each side by 6}

Sides of second triangle are 8, 16, and 24

For more help from me, visit: www.algebrahouse.com




Answer by stanbon(75887)

Two triangles are similar. The sides of the first triangle are 2, 4, and 6 units. The largest side of the second triangle is 24 units. Find the perimeter of the second triangle.
------
Similarity factor = 24/6 = 4
-----
Old perimeter:: 2+4+6 = 12
-----
New perimeter:: 4*12 = 48
---------------------
Cheers,
Stan H.
--------------